OpenStudy (anonymous):

for what value of k, the given expression is a perfect square k^2x^2+2(k+1)x+4

OpenStudy (anonymous):

please answer fast....its urgent...

OpenStudy (anonymous):

hint: use disc to solve it

OpenStudy (anonymous):

how??? 2nd term is not 4kx

OpenStudy (anonymous):

and u r supposed to find value of k

OpenStudy (kirbykirby):

ah yes I forgot the +1.

OpenStudy (kirbykirby):

you can solve for k by using the quadratic formula

OpenStudy (anonymous):

Determinant of quadratic must be equal to zero.

OpenStudy (anonymous):

determinant or discriminant??? if u r talking about discriminant then plz explain why discriminant is zero in this case......bcz according to my knowledge disc=0 iff roots are equal......plz explain if u can..........

OpenStudy (anonymous):

oh sorry discriminant :) :D

OpenStudy (anonymous):

roots are equal \(x_1=x_2=\frac{b}{a}\)\[k^2x^2+2(k+1)x+4=(ax-b)(ax-b)=(ax-b)^2\]which is a perfect square :)

OpenStudy (whpalmer4):

If it's a perfect square, the constant term of the binomial must be \(\sqrt{4} = 2\). Then our binomial is \((\sqrt{k^2x^2}+\sqrt{4}) = (kx+2)\) and when squared and expanded \[(kx+2)^2 = k^2x^2+4kx+4\]but we also have\[k^2x^2+2(k+1)x + 4\]so that means\[2(k+1)x = 4kx\]Discarding common factors\[k+1=2k\]\[k=1\] Solving via discriminant: discriminant = 0 if perfect square \[a = k^2\]\[ b=2(k+1)\]\[ c=4\] Discriminant \( = b^2-4ac = 0 = 2^2(k+1)^2-4*4*k^2\)\[(k+1)^2-4k^2=0\]\[k^2+2k+1-4k^2=0\]\[-3k^2+2k+1=0\]\[(3k+1)(1-k) = 0\]\[k=1,k=-1/3\]However \(k=-1/3\) is extraneous because we don't get a perfect square with it. Demonstration left for the reader :-)